# Discrete Chi-Square Method can model and forecast complex time series, like El Nino data between 1870 and 2024

### 04 Mar 2026 v7

@@ Acknowledgements
Figure S5: January data model $\mathcal{M}$=4 in Table S8. Red dots are -0.76 and -0.54 for 2025 and 2026. Otherwise as in Figure S2.
Table S8: Periods in January data. Note that we the pre-assigned sigficance level $\gamma=0.05$ (Equation 48) because monthly SST data are so noisy. Otherwise as in Table S1.
@@
Figure S6: February data model $\mathcal{M}$=4 in Table S9. Red dots are -0.35 and -0.40 for 2025 and 2026. Otherwise as in Figure S3
@@
$\mathcal{M}$=3 $4.7314PM0.050$ $5.672\pm 0.032$ $12.68\pm 0.44$ - - - $\uparrow$
$\mathcal{M}$3,1,1,R 11 $0.40\pm 0.11$ $0.450\pm 0.098$ $0.42\pm 0.10$ - - - 2.72 FebR14K311.dat
48.9 $1870.840\pm 0.69$ $1870.37\pm 0.54$ $1871.6\pm 1.6$ - - - 0.046
@@
$\mathcal{M}$=4 $3.771\pm 0.035$ $4.741\pm 0.058$ $5.647\pm 0.031$ $12.73\pm 0.22$ - - -
$\mathcal{M}$4,1,1,R 14 $0.3\pm 1.8$ $0.40\pm 0.10$ $0.45\pm 0.13$ $0.436\pm 0.094$ - - - FebR14K411.dat
46.2 $1873.8\pm 1.2$ $1870.8\pm 0.53$ $1870.68\pm 0.040$ $1871.32\pm 1.7$ - - -
@@
Figure S7: March data model $\mathcal{M}$=3 in Table S10. Red dot is -0.01 for 2025. Otherwise as in Figure S3
@@
$\mathcal{M}$=3 $3.633\pm 0.019$ $5.644\pm 0.052$ $12.73\pm 0.44$ - - - $\leftarrow$
$\mathcal{M}$3,1,1,R 11 $0.418\pm 0.090$ $0.45\pm 0.10$ $0.43\pm 0.13$ - - - 2.36 MarR14K311.dat
37.4 $1872.96\pm 0.37$ $1870.74\pm 0.81$ $1871.2\pm 1.2$ - - - 0.074
@@
Figure S8: April data model $\mathcal{M}$=4 in Table S11. Red dot is -0.20 for 2025. Otherwise as in Figure S3
@@
$\mathcal{M}$=1 $12.68\pm 0.13$ - - - $\uparrow$ $\uparrow$ $\uparrow$
@@
$\mathcal{M}$=3 $5.665\pm 0.017$ $12.73\pm 0.13$ $18.84\pm 0.84$ - - - $\uparrow$
$\mathcal{M}$3,1,1,R 11 $0.382\pm 0.097$ $0.416\pm 0.068$ $0.308\pm 0.065$ - - - 2.82 AprR14K311.dat
26.1 $1870.42\pm 0.29$ $1871.36\pm 0.85$ $1873.7\pm 2.0$ - - - 0.040
@@
Figure S9: May data model $\mathcal{M}$=4 in Table S12. Red dot is -0.04 for 2025. Otherwise as in Figure S3
@@
Figure S10: June data model $\mathcal{M}$=4 in Table S13. Red dot is -0.02 for 2025. Otherwise as in Figure S3
@@
$\mathcal{M}$=4 $3.647\pm 0.012$ $4.741PM0.037$ $12.76\pm 0.12$ $22.49\pm 1.7$ - - -
$\mathcal{M}$4,1,1,R 14 $0.289\pm 0.069$ $0.338\pm 0.094$ $0.415\pm 0.077$ $0.302\pm 0.086$ - - - JunR14K411.dat
21.8 $1872.07\pm 0.35$ $1870.23\pm 0.65$ $1871.35\pm 0.85$ $1891.0\pm 1.8$ - - -
@@
Figure S11: July data model $\mathcal{M}$=4 in Table S14. Red dot is -0.10 for 2025. Otherwise as in Figure S3
@@
Figure S13: September data model $\mathcal{M}$=3 in Table S16. Red dot is -0.40 for 2025. Otherwise as in Figure S3
@@
$\mathcal{M}$=2 $3.546\pm 0.022$ $20.10\pm 0.68$ - - - $\uparrow$ $\uparrow$
$\mathcal{M}$2,1,1,R 8 $0.349\pm 0.087$ $0.3550\pm 0.076$ - - - 2.70 2.66 Sep14K211.dat
41.1 $1871.60\pm 0.43$ $1871.6\pm 2.6$ - - - 0.048 0.018
@@
Figure S14: October data model $\mathcal{M}$=4 in Table S17. Red dot is -0.48 for 2025. Otherwise as in Figure S3
@@
Figure S15: Novermber data model $\mathcal{M}$=3 in Table S18. Red dot is -0.80 for 2025. Otherwise as in Figure S3
@@
Figure S16: December data model $\mathcal{M}$=4 in Table S19. Red dot is -0.92 for 2025. Otherwise as in Figure S3
@@
$\mathcal{M}$=1 $5.645\pm 0.033$ - - - $\uparrow$ $\uparrow$ $\uparrow$
$\mathcal{M}$1,1,1,R 5 $0.48\pm 0.13$ - - - 4.97 4.20 3.94 DecR14K111.dat
64.8 $1871.70\pm 0.44$ - - - 0.0026 0.00064 0.00017
@@
51.7 $1871.76\pm 0.38$ $1874.40\pm 0.38$ $1875.26\pm 0.50$ $1870.4\pm 2.0$ - - -
Table S20: “Big Wave” monthly forecast 2025-2036.
@@
2025 $0.3\pm 0.2$ $0.4\pm 0.2$ $0.2\pm 0.2$ $0.1\pm 0.2$ $0.2\pm 0.2$ $0.3\pm 0.2$ $2\pm 0.2 0.0$$0\pm 0.2$ $2\pm 0.2 0.0$$\pm 0.3 0.0$$0\pm 0.2$ $0.3\pm 0.2$
2026 $0.0\pm 0.3$ $1\pm 0.2 0.0$$0\pm 0.2$ $0.1\pm 0.1$ $0.3\pm 0.1$ $0.3\pm 0.2$ $0.1\pm 0.2$ $\pm 0.2 0.2±0.2$$\pm 0.2$ $0.2\pm 0.2$ $0.3\pm 0.3$ $0.2\pm 0.3$
2027 $0.3\pm 0.2$ $0.2\pm 0.3$ $0.4\pm 0.2$ $0.1\pm 0.2$ $0.4\pm 0.1$ $0.1\pm 0.1$ $0.1\pm 0.2$ $0.4\pm 0.1$ $0.2\pm 0.2$ $0.3\pm 0.2$ $0.3\pm 0.2$ $0.5\pm 0.2$
2028 $0\pm 0.2 0.0$$0\pm 0.2$ $\pm 0.2 0.2±0.2$$\pm 0.2$ $0.0\pm 0.2$ $0.1\pm 0.2$ $0.1\pm 0.2$ $0.2\pm 0.2$ $\pm 0.2 0.3±0.3$$\pm 0.3$ $1\pm 0.2 0.0$$0\pm 0.3$
2029 $0.3\pm 0.3$ $0.3\pm 0.2$ $0.2\pm 0.2$ $1\pm 0.1 0.0$$0\pm 0.2$ $0.2\pm 0.2$ $0.5\pm 0.2$ $1\pm 0.2 0.0$$0\pm 0.2$ $3\pm 0.2 0.0$$0\pm 0.3$ $0.1\pm 0.2$
2030 $0.2\pm 0.2$ $0.3\pm 0.2$ $\pm 0.2 0.1±0.1$$\pm 0.1$ $0.5\pm 0.2$ $0.4\pm 0.2$ $0.3\pm 0.2$ $4\pm 0.2 0.0$$0\pm 0.2$ $0.1\pm 0.2$ $0.6\pm 0.2$ $0.4\pm 0.2$
2031 $0.3\pm 0.3$ $0.3\pm 0.2$ $0.6\pm 0.2$ $0.6\pm 0.2$ $0.4\pm 0.2$ $0.3\pm 0.1$ $0.1\pm 0.1$ $0.5\pm 0.1$ $1\pm 0.2 0.0$$0\pm 0.2$ $0.7\pm 0.2$ $0.6\pm 0.2$
2032 $0.4\pm 0.2$ $0.4\pm 0.2$ $0.4\pm 0.2$ $0.7\pm 0.2$ $0.2\pm 0.1$ $1\pm 0.1 0.0$$0\pm 0.2$ $1\pm 0.1 0.0$$0\pm 0.3$ $0.3\pm 0.3$ $0.2\pm 0.3$ $0.3\pm 0.2$
2033 $0.1\pm 0.3$ $2\pm 0.2 0.0$$0\pm 0.2$ $0.4\pm 0.2$ $0.4\pm 0.2$ $0.3\pm 0.2$ $3\pm 0.1 0.0$$0\pm 0.2$ $0.2\pm 0.2$ $2\pm 0.3 0.0$$\pm 0.3 0.0$$0\pm 0.2$
2034 $0.2\pm 0.3$ $2\pm 0.3 0.0$$\pm 0.2 0.0$$0\pm 0.1$ $0.2\pm 0.2$ $0.5\pm 0.2$ $0.1\pm 0.2$ $1\pm 0.2 0.0$$0\pm 0.2$ $0.4\pm 0.3$ $0.1\pm 0.3$ $0.2\pm 0.3$
2035 $0.1\pm 0.3$ $0.1\pm 0.3$ $\pm 0.3 0.1±0.1$$\pm 0.1$ $0.4\pm 0.2$ $0.3\pm 0.2$ $\pm 0.1 0.2±0.2$-$\pm 0.2$ $0.2\pm 0.3$ $0.5\pm 0.3$ $0.1\pm 0.2$ $0.4\pm 0.3$
2036 $0\pm 0.3 0.0$$0\pm 0.2$ $0.3\pm 0.3$ $0.1\pm 0.2$ $0.1\pm 0.2$ $0.2\pm 0.1$ $0.2\pm 0.1$ $5\pm 0.1 0.0$$\pm 0.2 0.0$$0\pm 0.2$ $0.1\pm 0.2$ $0.1\pm 0.2$

### 24 Mar 2026 v8

@@ Acknowledgements
Figure S6: January data model $\mathcal{M}$=4 in Table S11. Red dots from Table S7 (Column 3). Otherwise as in Figure S2.
Table S11: Periods in January data. Note that we use the pre-assigned sigficance level $\gamma=0.05$ (Equation 48) because data of separate months data are so noisy. Otherwise as in Table S1.
@@
Figure S7: February data model $\mathcal{M}$=4 in Table S12. Otherwise as in Figure S6.
@@
$\mathcal{M}$=3 $4.741\pm 0.052$ $5.644\pm 0.089$ $12.73\pm 0.69$ - - - $\uparrow$
$\mathcal{M}$3,1,1,R 11 $0.40\pm 0.11$ $0.45\pm 0.11$ $0.43\pm 0.101$ - - - 2.72 FebR14K311.dat
48.9 $1870.960\pm 0.60$ $1870.88\pm 0.48$ $1871.37\pm 1.8$ - - - 0.046
@@
$\mathcal{M}$=4 $3.770\pm 0.034$ $4.742\pm 0.046$ $5.647\pm 0.031$ $12.73\pm 0.22$ - - -
$\mathcal{M}$4,1,1,R 14 $0.37\pm 0.11$ $0.398\pm 0.097$ $0.45\pm 0.13$ $0.436\pm 0.094$ - - - FebR14K411.dat
46.2 $1872.12\pm 0.44$ $1870.93\pm 0.67$ $1870.80\pm 0.47$ $1871.31\pm 1.5$ - - -
@@
Figure S8: March data model $\mathcal{M}$=3 in Table S13. Otherwise as in Figure S6.
@@
$\mathcal{M}$=3 $3.633\pm 0.019$ $5.671\pm 0.0289$ $12.68\pm 0.13$ - - - $\leftarrow$
$\mathcal{M}$3,1,1,R 11 $0.42\pm 0.12$ $0.45\pm 0.10$ $0.418\pm 0.090$ - - - 2.36 MarR14K311.dat
37.4 $1873.16\pm 0.26$ $1870.58\pm 0.50$ $1871.85\pm 0.77$ - - - 0.074
@@
Figure S9: April data model $\mathcal{M}$=4 in Table S14. Otherwise as in Figure S6.
@@
$\mathcal{M}$=1 $12.69\pm 0.13$ - - - $\uparrow$ $\uparrow$ $\uparrow$
@@
$\mathcal{M}$=3 $5.665\pm 0.017$ $12.73\pm 0.13$ $18.83\pm 0.84$ - - - $\uparrow$
$\mathcal{M}$3,1,1,R 11 $0.382\pm 0.097$ $0.416\pm 0.068$ $0.31\pm 0.11$ - - - 2.82 AprR14K311.dat
26.1 $1870.71\pm 0.48$ $1871.65\pm 0.89$ $1874.0\pm 1.6$ - - - 0.040
@@
Figure S10: May data model $\mathcal{M}$=4 in Table S15. Otherwise as in Figure S6.
@@
Figure S11: June data model $\mathcal{M}$=4 in Table S16. Otherwise as in Figure S6.
@@
$\mathcal{M}$=4 $3.647\pm 0.012$ $4.741\pm 0.037$ $12.76\pm 0.12$ $22.44\pm 1.7$ - - -
$\mathcal{M}$4,1,1,R 14 $0.289\pm 0.069$ $0.338\pm 0.09$ $0.415\pm 0.077$ $0.302\pm 0.086$ - - - JunR14K411.dat
21.8 $1872.53\pm 0.35$ $1870.8\pm 0.36$ $1871.81\pm 0.99$ $1891.5\pm 1.8$ - - -
@@
Figure S12: July data model $\mathcal{M}$=4 in Table S17. Otherwise as in Figure S6.
@@
Figure S14: September data model $\mathcal{M}$=3 in Table S19. Otherwise as in Figure S6.
@@
$\mathcal{M}$=2 $3.546\pm 0.022$ $20.1\pm 1.3$ - - - $\uparrow$ $\uparrow$
$\mathcal{M}$2,1,1,R 8 $0.35\pm 0.13$ $0.350\pm 0.076$ - - - 2.70 2.66 Sep14K211.dat
41.1 $1872.32\pm 0.34$ $1872.3\pm 2.9$ - - - 0.048 0.018
@@
Figure S15: October data model $\mathcal{M}$=4 in Table S20. Otherwise as in Figure S6.
@@
Figure S16: November data model $\mathcal{M}$=3 in Table S21. Otherwise as in Figure S6.
@@
Figure S17: December data model $\mathcal{M}$=4 in Table S22. Otherwise as in Figure S6.
@@
$\mathcal{M}$=1 $3.645\pm 0.019$ - - - $\uparrow$ $\uparrow$ $\uparrow$
$\mathcal{M}$1,1,1,R 5 $0.48\pm 0.12$ - - - 4.97 4.20 3.94 DecR14K111.dat
64.8 $1872.66\pm 0.45$ - - - 0.0026 0.00064 0.00017
@@
51.7 $1872.71\pm 0.38$ $1875.35\pm 0.59$ $1876.22\pm 0.80$ $1871.4\pm 1.2$ - - -
Table S23: Separate months. “Big Wave” forecast (Figures S6-S17: open circles). Otherwise as in Table S10.
@@
2025 $0.3\pm 0.3$ $\uparrow$ $0.4\pm 0.3$ $\uparrow$ $0.2\pm 0.2 \downarrow$ $0.1\pm 0.2 \downarrow$ $0.2\pm 0.2 \downarrow$ $0.3\pm 0.1 \downarrow$ $2\pm 0.2 \uparrow 0.0$$0\pm 0.2 \downarrow$ $2\pm 0.2 \uparrow 0.0$$\pm 0.2 \downarrow 0.0$$0\pm 0.2 \downarrow$ $0.3\pm 0.2 \downarrow$
2026 $0.0\pm 0.2 \uparrow$ $1\pm 0.2 \uparrow 0.0$$0\pm 0.2 \uparrow$ $0.1\pm 0.1 \downarrow$ $0.3\pm 0.2$ $\uparrow$ $0.3\pm 0.1 \downarrow$ $0.1\pm 0.2$ $\uparrow$ $\pm 0.2 \uparrow 0.2±0.2$$\pm 0.2$ $\uparrow$ $0.2\pm 0.2$ $\uparrow$ $0.3\pm 0.2$ $\uparrow$ $0.2\pm 0.2$ $\downarrow$
2027 $0.3\pm 0.3 \downarrow$ $0.2\pm 0.2 \downarrow$ $0.4\pm 0.2$ $\uparrow$ $0.1\pm 0.2$ $\uparrow$ $0.4\pm 0.2$ $\uparrow$ $0.1\pm 0.1 \downarrow$ $0.1\pm 0.2 \uparrow$ $0.4\pm 0.2$ $\uparrow$ $0.2\pm 0.2 \downarrow$ $0.3\pm 0.3$ $\uparrow$ $0.3\pm 0.2$ $\uparrow$ $0.5\pm 0.2$ $\uparrow$
2028 $0.1\pm 0.4$ $\downarrow$ $0.1\pm 0.2$ $\downarrow$ $0.2\pm 0.2$ $\uparrow$ $2\pm 0.1 \uparrow 0.0$$0\pm 0.2 \downarrow$ $\pm 0.2 \uparrow 0.1±0.1$$\pm 0.1$ $\uparrow$ $0.2\pm 0.2$ $\uparrow$ $0.3\pm 0.2 \downarrow$ $0.3\pm 0.2$ $\uparrow$ $1\pm 0.3 \downarrow 0.0$$0\pm 0.2 \uparrow$
2029 $0.3\pm 0.2$ $\uparrow$ $0.3\pm 0.2$ $\uparrow$ $0.2\pm 0.2 \downarrow$ $1\pm 0.2 \downarrow 0.0$$0\pm 0.2 \downarrow$ $0.2\pm 0.2$ $\uparrow$ $0.5\pm 0.1$ $\uparrow$ $1\pm 0.1 \downarrow 0.0$$0\pm 0.2 \downarrow$ $3\pm 0.3 \uparrow 0.0$$0\pm 0.2 \downarrow$ $0.1\pm 0.3 \downarrow$
2030 $0.2\pm 0.3$ $\uparrow$ $0.3\pm 0.2$ $\uparrow$ $0.1\pm 0.2$ $\downarrow$ $0.1\pm 0.2$ $\downarrow$ $0.5\pm 0.2$ $\uparrow$ $0.4\pm 0.1$ $\downarrow$ $0.3\pm 0.2$ $\downarrow$ $4\pm 0.2 \uparrow 0.0$$0\pm 0.2 \downarrow$ $0.1\pm 0.2$ $\uparrow$ $0.6\pm 0.2$ $\uparrow$ $0.4\pm 0.2$ $\downarrow$
2031 $0.3\pm 0.3$ $\downarrow$ $0.4\pm 0.2$ $\downarrow$ $0.6\pm 0.1$ $\uparrow$ $0.6\pm 0.2$ $\uparrow$ $0.4\pm 0.1$ $\downarrow$ $0.3\pm 0.1$ $\downarrow$ $0.1\pm 0.1 \downarrow$ $0.5\pm 0.2$ $\uparrow$ $1\pm 0.2 \downarrow 0.0$$0\pm 0.2 \uparrow$ $0.7\pm 0.1$ $\uparrow$ $0.6\pm 0.3$ $\downarrow$
2032 $0.4\pm 0.2$ $\downarrow$ $0.4\pm 0.2$ $\uparrow$ $0.4\pm 0.2$ $\uparrow$ $0.7\pm 0.2$ $\uparrow$ $0.2\pm 0.2$ $\downarrow$ $1\pm 0.2 \downarrow 0.0$$0\pm 0.1 \downarrow$ $1\pm 0.2 \uparrow 0.0$$0\pm 0.2 \downarrow$ $0.3\pm 0.2$ $\uparrow$ $0.2\pm 0.2$ $\downarrow$ $0.3\pm 0.2$ $\uparrow$
2033 $0.1\pm 0.3$ $\downarrow$ $1\pm 0.2 \uparrow 0.0$$0\pm 0.2 \downarrow$ $0.4\pm 0.2$ $\uparrow$ $0.4\pm 0.2$ $\downarrow$ $0.3\pm 0.2$ $\downarrow$ $3\pm 0.1 \downarrow 0.0$$0\pm 0.2 \downarrow$ $0.2\pm 0.2$ $\uparrow$ $2\pm 0.3 \downarrow 0.0$$\pm 0.3 \downarrow 0.0$$0\pm 0.3 \downarrow$
2034 $\pm 0.3 \downarrow 0.2±0.2$-$\pm 0.2$ $\downarrow$ $0\pm 0.2 \uparrow 0.0$$0\pm 0.2 \uparrow$ $0.2\pm 0.2$ $\uparrow$ $0.5\pm 0.2$ $\uparrow$ $0.1\pm 0.2$ $\downarrow$ $1\pm 0.2 \uparrow 0.0$$0\pm 0.2 \downarrow$ $0.4\pm 0.3 \downarrow$ $0.1\pm 0.2$ $\uparrow$ $0.2\pm 0.2 \downarrow$
2035 $0.1\pm 0.3 \uparrow$ $0.1\pm 0.2 \uparrow$ $0.1\pm 0.2 \downarrow$ $0.1\pm 0.2$ $\uparrow$ $0.4\pm 0.2 \downarrow$ $0.3\pm 0.2$ $\uparrow$ $\pm 0.2 \downarrow 0.2±0.2$-$\pm 0.2$ $\uparrow$ $0.2\pm 0.2 \uparrow$ $0.5\pm 0.2 \downarrow$ $0.1\pm 0.2 \uparrow$ $0.4\pm 0.2 \downarrow$
2036 $0\pm 0.3 \uparrow 0.0$$0\pm 0.3 \uparrow$ $0.3\pm 0.2 \downarrow$ $0.1\pm 0.2$ $\uparrow$ $0.1\pm 0.2 \downarrow$ $\pm 0.2 \downarrow 0.2±0.2$-$\pm 0.2$ $\uparrow$ $5\pm 0.1 \downarrow 0.0$$\pm 0.2 \uparrow 0.0$$0\pm 0.3 \uparrow$ $0.1\pm 0.2 \downarrow$ $0.1\pm 0.2 \uparrow$
[…]
Line $P_{1}$ $P_{2}=10$ $P_{2}=11$ $P_{2}=11.86$
(-) (y) (y) (y) (y)
(1) (2) (3) (4) (5)
(1) 10 - 110 63.8
(2) - 5.23 5.42
(3) 10/2 - 9.16 8.64
(4) - 3.43 3.51
(5) 10/3 - 4.78 4.63
(6) - 2.55 2.60
(7) 11 110 - 152
(8) 5.23 - 5.70
(9) 11/2 12.22 - 10.25
(10) 3.55 - 3.75
(11) 11/3 5.79 - 5.31
(12) 2.68 - 2.80
(13) 11.86 63.8 152 -
(14) 5.42 5.71 -
(15) 11.86/2 14.57 12.86 -
(16) 3.72 3.85 -
(17) 11.86/3 6.53 6.17 -
(18) 2.83 2.90 -